### v1

# Collaborative Perception Datasets for Autonomous Driving: A Review

Abstract
Collaborative perception has attracted growing interest from academia and industry due to its potential to enhance perception accuracy, safety, and robustness in autonomous driving through multi-agent information fusion. With the advancement of Vehicle-to-Everything (V2X) communication, numerous collaborative perception datasets have emerged, varying in cooperation paradigms, sensor configurations, data sources, and application scenarios. However, the absence of systematic summarization and comparative analysis hinders effective resource utilization and standardization of model evaluation. As the first comprehensive review focused on collaborative perception datasets, this work reviews and compares existing resources from a multi-dimensional perspective. We categorize datasets based on cooperation paradigms, examine their data sources and scenarios, and analyze sensor modalities and supported tasks. A detailed comparative analysis is conducted across multiple dimensions. We also outline key challenges and future directions, including dataset scalability, diversity, domain adaptation, standardization, privacy, and the integration of large language models. To support ongoing research, we provide a continuously updated online repository of collaborative perception datasets and related literature: https://github.com/frankwnb/Collaborative-Perception-Datasets-for-Autonomous-Driving.
Index Terms:
Autonomous Driving, Collaborative Perception, Datasets, V2X Communication, Perception Tasks
I Introduction
Perception serves as a foundational component of Autonomous Driving (AD) systems [1], [2], [3], [4], [5], [6], [7], [8], [9], enabling the vehicle to perceive its surroundings through various onboard sensors and to perform real-time tasks such as object detection [10][11], semantic segmentation, tracking, and trajectory prediction in dynamic traffic environments [12]. The precision and reliability of the perception system directly impact the safety and decision-making capability of autonomous vehicles, making it a critical aspect to achieve high-level AD. In recent years, significant progress has been made in single-vehicle perception, driven by progress in computer vision [13][14][15], deep learning [16][17], and sensor technology [18][19]. Numerous studies have leveraged cameras [20][21][22], LiDAR [23][24], Radar [25][26], and multi-sensor fusion [27][28][29][30][31][32] to enhance environmental perception in AD systems. However, traditional single-vehicle perception is constrained by limited fields of view, sensor blind spots, and environmental occlusions, making it insufficient to meet the high-reliability requirements of AD in complex and dynamic traffic environments.
Figure 1: Development Trends of Collaborative Perception Datasets.
To overcome the limitations of single-vehicle perception, Collaborative Perception (CP) leverages Vehicle-to-Everything (V2X) communication to enable real-time information sharing among vehicles and infrastructure [33], thereby enhancing the perception of occluded and long-range regions, improving driving safety, and expanding the Operational Design Domain (ODD) [34][35]. For instance, at an urban intersection, a single vehicle may fail to detect pedestrians or oncoming vehicles that are occluded by buildings or other obstacles. By leveraging V2X connectivity, autonomous vehicles can receive supplementary perception data from roadside infrastructure or neighboring vehicles, thereby improving situational awareness and reducing accident risk. V2X communication encompasses Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), Infrastructure-to-Infrastructure (I2I), and Vehicle-to-Network (V2N) links, empowering AD systems to overcome the inherent limitations of individual perception and achieve a more holistic understanding of their surroundings.
High-quality datasets are essential for advancing algorithm development and performance evaluation in CP research [36]. Existing single-vehicle perception datasets, such as KITTI [37], nuScenes [38], and Waymo [2], are primarily designed for individual perception tasks and do not support multi-vehicle or infrastructure-level information exchange, rendering them unsuitable for direct use in CP studies.
In recent years, numerous CP datasets, as illustrated in Fig. 1, have been released to support various scenarios, tasks, and sensor modalities, including DAIR-V2X [39], OPV2V [40], V2X-Sim [41], and V2X-Seq [42]. While these datasets offer valuable resources for the CP community, they differ significantly in terms of scale, scene diversity, annotation quality, sensor configurations, and the simulation of communication latency. A systematic review and analysis of existing CP datasets can help researchers better understand their characteristics, application boundaries, and limitations, thereby providing scientific guidance for future investigations. Furthermore, identifying existing gaps and prospective research directions will contribute to the continued advancement of CP technologies.
I-A Roadmap
Figure 2: Chronological overview of collaborative perception datasets, including roadside perception, V2V, V2I, V2X, and I2I datasets.
Fig. 2 illustrates the chronological evolution of CP datasets. Overall, the development of these datasets has progressed through several key phases, each marking a significant shift in perception paradigms or improvements in dataset diversity. Changes in collaboration models, increasing scene complexity, the expansion of supported tasks, and the continuous growth in sensor configurations have all contributed to the advancement of CP technologies.
In the early stages, datasets were primarily focused on infrastructure-based perception, as exemplified by Ko-PER [43], CityFlow [44], and the INTERACTION [45]. However, these datasets did not support V2V or V2I information exchange, thereby limiting their applicability to CP research. With the advancement of V2X communication technologies, the number of related datasets has steadily increased. Around 2020, several V2V and V2I datasets were released, including T & J [46], V2V-Sim [47], CoopInf [48], and CODD [49], which collectively accelerated progress in CP research. However, datasets from this period still faced limitations in terms of scene complexity, sensor modality diversity, and task coverage. As the field progressed, newer CP datasets began to target a broader range of scenarios. For example, Adver-City [50] focuses on adverse weather conditions, DeepAccident [51] addresses accident prediction, and H-V2X [52] is tailored for highway environments. These datasets have significantly enhanced the adaptability of CP systems.
Since 2022, CP dataset development has increasingly emphasized multi-modal fusion and real-world deployment. Datasets such as V2V4Real [53], V2X-Real [54], and DAIR-V2X-C [39] incorporate real-world data collection to overcome the limitations of synthetic datasets in terms of authenticity and distribution. Simultaneously, newer datasets demonstrate broader task coverage and richer sensor modalities. For instance, V2X-Seq [42], V2XReID [55], and DAIR-V2X-Traj [56] support a variety of perception tasks, while V2X-Radar [57] and V2X-R [58] incorporate 4D radar to enhance perception quality. Moreover, the emergence of I2I datasets such as InScope [59] and Rcooper [60] has further enriched the collaborative paradigms within CP. In general, the evolution of CP datasets reflects a continuous trend toward greater complexity, richer modalities, and more specialized scenarios.
I-B Comparison to Related Reviews
We aim to clarify the distinctions between this review and previous related reviews [61], [36], [62], [63], [64], [65], [66], [67], [68], [69], [70]. While some existing reviews partially overlap with ours in terms of dataset coverage [71] [72], they omit several important CP datasets and recent benchmark methodologies introduced in the past few years. Furthermore, their classification frameworks tend to be narrow and often lack systematic summarization and in-depth comparative analysis. In addition, some reviews concentrate primarily on vehicle-centric datasets [61][36][62][63], or explore specific aspects of CP such as multimodal fusion [64][65] or certain CP techniques [66][67][68][69][70], without offering a comprehensive overview of dataset evolution and their applicability across various scenarios. In contrast, this review presents the most up-to-date and complete collection of CP datasets, encompassing a broader spectrum of perspectives. It also delivers an in-depth discussion on dataset diversity, key challenges, and emerging research directions.
Figure 3: Organization of this paper.
I-C Main Contributions
In this review, we present a systematic and comprehensive analysis of CP datasets for AD, as summarized in Fig. 3. The main contributions of this work are as follows:
Comprehensive Review of CP Datasets: To the best of our knowledge, this is the first dedicated and comprehensive review focused specifically on CP datasets. We systematically categorize and analyze available datasets based on collaboration types, data sources, sensor modalities, and application scenarios.
Task-Oriented Dataset Classification: We classify CP datasets according to core perception tasks in AD, including 3D object detection, semantic segmentation, and object tracking and prediction. This task-centric taxonomy facilitates effective dataset selection based on research needs. In addition, we cover the latest developments in CP datasets, providing a timely and up-to-date overview.
Dataset Comparison and Emerging Trends: We conduct multi-dimensional comparisons, highlight high-impact datasets, and discuss key challenges and future directions such as standardization, LLM-driven annotation, and data diversity.
I-D Review Structure
The structure of this paper is illustrated in Fig. 3. To better present our review, the remainder of this paper is organized as follows. Section II introduces datasets under different collaborative paradigms. Section III compares the characteristics of simulated and real-world datasets. Section IV reviews existing CP datasets from the perspective of different application scenarios. Section V analyzes datasets based on sensor modalities. Section VI summarizes and analyzes the CP tasks supported by these datasets, along with their corresponding evaluation metrics. Section VII provides a comprehensive analysis of existing CP datasets, focusing on their availability, quality, and impact. Section VIII discusses the challenges and trends in CP datasets. Section IX concludes the paper and outlines future research directions. A summary of all abbreviations used throughout the paper is provided in Table I.
Table I: Summary of commonly used abbreviations in this review.
Abbreviation Definition
AD Autonomous Driving
AV Autonomous Vehicle
CAV Connected and Autonomous Vehicle
VIC vehicle-infrastructure cooperation
CP Cooperative or Collaborative Perception
V2X Vehicle-to-Everything
V2V Vehicle-to-Vehicle
V2I Vehicle-to-Infrastructure
V2N Vehicle-to-Network
I2I Infrastructure-to-Infrastructure
RSU Roadside Unit
ITS Intelligent Transportation System
CAM Cooperative Awareness Message
CPM Cooperative Perception Message
CPS Collective Perception Service
CAD Collaborative Autonomous Driving
CDA Cooperative Driving Automation
GPS Global Positioning System
IMU Inertial Measurement Unit
BEV Bird’s Eye View
CNN Convolutional Neural Network
MF Motion Forecasting
Re-ID Re-Identification
LLM Large Language Model
II Collaboration-Based Perception Datasets
With the rapid development of Intelligent Connected Vehicles (ICVs) and Intelligent Transportation Systems (ITS), CP paradigms have become increasingly diverse. These paradigms now encompass not only pure roadside perception but also V2V, V2I, and more comprehensive V2X. Additionally, emerging paradigms such as I2I collaboration have recently gained attention. As illustrated in Fig. 4, various CP architectures are now being explored. In this section, we classify and discuss CP datasets according to their communication paradigms, including roadside perception datasets (Section II-A), V2V datasets (Section II-B), V2I datasets (Section II-C), V2X datasets (Section II-D), and I2I datasets (Section II-E). Table II provides a comprehensive summary of representative datasets under each paradigm. These datasets play a crucial role in advancing cooperative intelligence for ICVs and ITS.
Table II: Publicly Available Datasets for Collaborative Perception
View Dataset Year Venue Sensors Source Tasks Size Link
Image LiDAR 3D Box
Infra Ko-PER [43] 2014 ITSC C, L Real 3DOD, MOT 18.7k 4.8k - Link
CityFlow [44] 2019 CVPR C Real MTSCT/MTMCT, ReID 118k - 22.9k Link
INTERACTION [45] 2019 IROS C, L Real 2DOD, TP 1.4M - 1.4M Link
A9-Dataset [73] 2022 IV C, L Real 3DOD 1.1k - 14.4k Link
IPS300+ [74] 2022 ICRA C, L Real 2DOD, 3DOD 56.7k 14.2k 4.5M Link
Rope3D [75] 2022 CVPR C, L Real 2DOD, 3DOD 50k - 1.5M Link
LUMPI [76] 2022 IV C, L Real 3DOD 200k 90k - Link
TUMTraf-I [77] 2023 ITSC C, L Real 3DOD 4.8k 4.8k 57.4k Link
RoScenes [78] 2024 ECCV C Real 3DOD 1.3M - 21.13M Link
H-V2X [52] 2024 ECCV C, R Real BEV Det, MOT, TP 1.94M - - Link
V2V COMAP [79] 2021 ISPRS L, C Sim 3DOD, SS 8.6k - 226.9k Link
CODD [49] 2021 RA-L L Sim Registration - 13.5k 204k Link
OPV2V [40] 2022 ICRA C, L, R Sim 3DOD 44k 11.4k 232.9k Link
OPV2V+ [80] 2023 CVPR C, L, R Sim 3DOD 11.4k+ - 232.9k+ Link
V2V4Real [53] 2023 CVPR L, C Real 3DOD, MOT, S2R 40k 20k 240k Link
LUCOOP [81] 2023 IV L Real 3DOD - 54k 7k Link
MARS [82] 2024 CVPR L, C Real VPR, NR 1.4M - - Link
OPV2V-H [83] 2024 ICLR C, L, R Sim 3DOD 79k - 232.9k+ Link
V2V-QA [84] 2025 arXiv L, C Real 3DOD, PQA - 18k - Link
CP-UAV [35] 2022 NIPS L, C Sim 3DOD 131.9K - 1.94M Link
V2I CoopInf [48] 2020 TITS L, C Sim 3DOD 10k - 121.2K Link
DAIR-V2X-C [39] 2022 CVPR L, C Real 3DOD 71k 71k 1.2M Link
V2X-Seq [42] 2023 CVPR L, C Real 3DOT, TP 71k 15k 464k Link
HoloVIC [85] 2024 CVPR L, C Real 3DOD, MOT 100k - 11.47M Link
OTVIC [86] 2024 IROS L, C Real 3DOD 15k - 24.4k Link
DAIR-V2XReid [55] 2024 TITS L, C Real 3DOD, Re-ID 7.3k - - Link
TUMTraf V2X [87] 2024 CVPR L, C Real 3DOD, MOT 5k 2k 30k Link
V2X-Radar [57] 2024 arxiv L, C, R Real 3DOD 40k 20k 350k Link
V2X V2X-Sim [41] 2022 RA-L L, C Sim 3DOD, MOT, SS 60k 10K 26.6K Link
V2XSet [34] 2022 ECCV L, C Sim 3DOD 44k 11.4k 233k Link
DOLPHINS [88] 2022 ACCV L, C Sim 2DOD, 3DOD 42.3k 42.3k 292.5k Link
DeepAccident [51] 2024 AAAI L, C Sim 3DOD, MOT, SS, TP - 57k 285k Link
V2X-Real [54] 2024 ECCV L, C Real 3DOD 171k 33k 1.2M Link
Multi-V2X [89] 2024 arxiv L, C Sim 3DOD, MOT 549k 146k 4.2M Link
Adver-City [50] 2024 arxiv L, C Sim 3DOD, MOT, SS 24k 24k 890k Link
DAIR-V2X-Traj [56] 2024 NIPS L, C Real MF 808k - 1.4M Link
WHALES [90] 2024 arxiv L, C Sim 3DOD 70k 17k 2.01M Link
V2X-R [58] 2024 arxiv L, C, R Sim 3DOD 150.9k 37.7k 170.8k Link
V2XPnP-Seq [91] 2024 arxiv L, C Real Perception and Prediction 208k 40k 1.45M Link
Mixed Signals [92] 2025 arxiv L Real 3DOD - 45.1k 240.6k Link
SCOPE [93] 2024 arxiv C, L Sim 2DOD, 3DOD, SS, S2R 17k 17k 575k Link
I2I Rcooper [60] 2024 CVPR C, L Real 3DOD, MOT 50k 30k 30k Link
InScope [59] 2024 arxiv L Real 3DOD, MOT - 21.3k 188k Link
Note: Sensors: Camera (C), LiDAR (L), Radar (R). Source: Real = collected in the real world; Sim = generated via simulation. Tasks: 2DOD = 2D Object Detection, 3DOD = 3D Object Detection, MOT = Multi-Object Tracking, MTSCT = Multi-target Single-camera Tracking, MTMCT = Multi-target Multi-camera Tracking, SS = Semantic Segmentation, TP = Trajectory Prediction, VPR = Visual Place Recognition, NR = Neural Reconstruction, Re-ID = Re-Identification, S2R = Sim2Real, MF = Motion Forecasting, PQA = Planning Q&A.
Figure 4: Schematic diagram of different collaborative perception methods, including roadside perception, V2V, V2I, V2X, and I2I.
II-A Roadside Datasets
Roadside perception plays a vital role in AD and ITS by offering a global, occlusion-resilient view of traffic participants [94][95][96]. Unlike vehicle-mounted systems, infrastructure-based sensing units—typically installed on poles, gantries, or overpasses—enable the monitoring of complex intersections, dense traffic flows, and long-range environments with enhanced spatial stability and temporal continuity. By supplementing on-vehicle sensing with infrastructure-level observations, roadside perception significantly improves perception robustness in scenarios involving occlusions, blind spots, or congested conditions.
A variety of representative roadside datasets have been introduced to support infrastructure-based perception research. Early datasets such as Ko-PER [43] and CityFlow [44] primarily target intersection environments and person or vehicle Re-ID tasks. INTERACTION [45] expands the behavioral diversity of roadside scenarios by incorporating multi-national data with adversarial and cooperative driving maneuvers. Rope3D [75] provides high-resolution 3D data along with dedicated metrics for evaluating roadside 3D object detection. LUMPI [76] supports multi-view analysis using synchronized multi-sensor setups, while IPS300+ [74] offers dense, long-range 3D annotations for complex urban intersections. TUMTraf-I [77] enhances diversity through coverage of varying lighting and weather conditions. For highway-oriented research, A9-Dataset [73] offers real-world, high-speed traffic scenarios, whereas RoScenes [78] features the largest perception area to date through high-density, multi-camera roadside deployment. H-V2X [52] extends this trend by integrating radar and camera data, supporting a range of tasks including detection, tracking, and trajectory prediction under complex highway conditions.
The release of roadside datasets has significantly broadened the scope of perception research, enabling detailed studies on object detection, semantic segmentation, and multi-object tracking from an infrastructure-centric perspective. These datasets also provide a valuable foundation for future V2I CP systems by enhancing scene coverage and perceptual redundancy. However, several challenges remain inherent to infrastructure-based data collection. Such as, high deployment and maintenance costs restrict large-scale installations, particularly in highway and rural areas. Data distributions are often skewed toward urban environments, which limits the generalizability of trained models across diverse traffic scenarios. Additionally, precise temporal synchronization and spatial alignment among heterogeneous sensors remain non-trivial, especially in long-term or multi-site deployments.
II-B V2V Datasets
V2V CP enhances perception range and accuracy by enabling information sharing among multiple vehicles. Compared to single-vehicle perception, V2V collaboration effectively mitigates blind spots caused by occlusions and compensates for the limitations of individual sensors in specific scenarios. This leads to improved robustness in tasks such as object detection, semantic understanding, and behavior prediction. Furthermore, the integration of multi-sensor data and cross-view information enables V2V systems to perform more adaptively in complex and dynamic environments. Therefore, the development of high-quality V2V CP datasets is of great importance for advancing research in multi-vehicle collaborative autonomous driving systems.
In recent years, V2V datasets have continued to evolve in terms of collaborative modeling and data organization mechanisms. Most datasets, including OPV2V [40], MARS [82], and LUCOOP [81], adopt an offline alignment strategy that unifies data from multiple vehicles into a common coordinate system using global positioning or timestamps. This approach simulates idealized synchronized collaboration scenarios. However, it fails to account for real-world issues such as asynchronous sampling, communication latency, and localization errors during actual V2V communication. To address these limitations, HEAL [83] introduces asynchronous perception and temporal alignment mechanisms, while OPV2V-N [97] employs neural modeling to improve perception robustness. COMAP [79] adopts a task-oriented modeling approach from a multi-agent SLAM perspective. In terms of agent configuration, most datasets support CP among 2 to 4 vehicles. MARS [82] enables data collection across multiple time periods, trajectories, and viewpoints using four connected vehicles. OPV2V+ [80] extends viewpoint variation modeling by supporting primary view selection and perception completion. Regarding task coverage, existing datasets support key tasks such as object detection, tracking, Bird’s Eye View (BEV) semantic segmentation, and path planning. V2V-QA [84] further incorporates multi-modal large language models to support intelligent question answering and path reasoning tasks. IRV2V [98] and CODD [49] focus on asynchronous robustness and multi-view point cloud registration, reflecting the expanding scope of V2V data modeling in terms of collaboration granularity and semantic capabilities.
Despite the notable progress made by V2V CP datasets in enabling multi-vehicle fusion research, several distinct challenges remain. For example, most existing datasets rely on offline stitching techniques for data integration and lack modeling of real communication links, transmission latency, and collaborative constraints. As a result, they fail to capture the uncertainty and dynamic characteristics of information exchange in real V2V scenarios, limiting deeper investigations into the essence of cooperative mechanisms. Looking ahead, the development of V2V datasets should focus on realistic modeling of communication processes, the design of dynamic alignment mechanisms for CP, and the construction of efficient multi-vehicle data acquisition and annotation pipelines in complex urban environments.
II-C V2I Datasets
V2I perception datasets aim to enhance AD systems by enabling collaboration between vehicles and roadside infrastructure. These datasets incorporate multimodal data from both vehicle-mounted and infrastructure-based sensors, offering a more comprehensive and accurate understanding of the environment.
Representative V2I datasets exhibit distinct characteristics and collaborative strategies. CoopInf [48] employs eight infrastructure sensors for multi-view 3D object detection. DAIR-V2X-Seq [42] is the first large-scale sequential V2X dataset using data from both infrastructure and vehicles. HoloVIC [85] constructs holographic intersections with diverse layouts to overcome occlusion and blind spots through multisensor fusion. TUMTraf-V2X [87] emphasizes perception under complex lighting conditions and supports 3D detection and tracking. OTVIC [86] targets high-speed and noisy communication scenarios with multimodal, multi-view perception data.
While these datasets significantly contribute to improving V2I CP, challenges remain. Sensor synchronization is technically demanding, and many datasets offer limited scene diversity or weather coverage. Future efforts should prioritize building more diverse, comprehensive V2I datasets to improve perception robustness in complex and dynamic driving environments.
II-D V2X Datasets
CP empowered by V2X technology facilitates data sharing and joint environmental understanding through communication among diverse agents, including vehicles, infrastructure, pedestrians, and other IoT devices. This paradigm significantly improves perception robustness and driving safety in autonomous systems.
From the perspective of agent structure and heterogeneity, recent V2X datasets reveal an increasing trend towards modeling complex interactions among heterogeneous agents with varying sensing capabilities. Early works such as V2X-SIM [41] support basic cooperative tasks, yet remain limited in agent type diversity and environmental complexity. The DAIR-V2X series [39, 55, 56] significantly expands the agent spectrum by introducing bi-agent collaboration between vehicles and RoadSide Units (RSUs). These datasets illustrate asymmetric collaboration in heterogeneous settings, where infrastructure and mobile vehicles possess distinct sensor setups and motion patterns.
To capture multi-agent temporal dynamics, V2XPnP [91] offers sequential data involving multiple cooperative vehicles, facilitating trajectory prediction in temporally correlated environments. DOLPHINS [88] further enhances agent diversity by simulating tri-vehicle collaboration in adverse conditions such as rain and fog, which supports evaluation of perception robustness under extreme scenarios. V2X-Real [54] emphasizes large-scale data collection across multiple vehicles and infrastructures, reflecting real-world V2X deployments with rich agent-level interactions. Adver-City [50] and DeepAccident [51] introduce environmental challenges such as snow, low visibility, and accident-prone zones, where CP across various agents becomes critical. Meanwhile, V2X-Radar [57] and V2X-R [58] expand the sensor modalities by incorporating 4D radar into the V2X sensing framework, highlighting the potential of cross-agent, cross-modal fusion in enhancing perception fidelity.
Despite these advancements, V2X datasets still face challenges related to scalability, diversity, and task integration. Many datasets lack coverage of complex urban intersections or multi-agent scenarios with mixed mobility patterns. Additionally, existing benchmarks often address specific tasks in isolation, limiting their support for evaluating end-to-end CP systems across heterogeneous agents.
II-E I2I Datasets
As V2X CP continues to evolve, I2I collaboration has emerged as a promising paradigm to address the limitations of single-agent roadside systems. While V2I, V2I and V2X datasets focus primarily on perception enhancement via VIC, their perception range is often constrained by occlusion, especially when large vehicles or urban structures obstruct the view. I2I CP addresses this issue by enabling multiple roadside infrastructure nodes to share and fuse data, thereby reducing blind spots and improving the robustness of perception in complex environments such as intersections and road corridors.
Two pioneering datasets exemplify the advancement of I2I CP. The RCooper [60] is the first real-world, large-scale roadside CP dataset. It includes two typical scenarios—intersection and corridor—with infrastructure agents equipped with multi-type LiDARs and cameras. RCooper provides high-quality annotations, demonstrating the effectiveness of cross-infrastructure cooperation. In contrast, the InScope dataset [59] specifically targets occlusion-aware perception by deploying LiDARs at multiple viewpoints. Additionally, InScope proposes a novel quantitative metric to evaluate anti-occlusion capabilities, providing a rigorous framework for I2I perception evaluation. V2XPnP [91] further expands the paradigm by integrating large-scale sequential data across V2X modes, particularly I2I, offering valuable support for complex CP tasks. Compared to other paradigms, I2I collaboration is more effective in mitigating occlusion in dense traffic environments. However, Current I2I datasets face challenges in generalization, data synchronization, and scene diversity, highlighting the need for scalable deployment strategies and standardized evaluation protocols.
III Simulation vs. Real-World Datasets
Simulation and real-world datasets serve complementary roles in AD research, differing in cost, source, annotation accuracy, realism, and applicability. Simulation datasets, generated through virtual modeling, offer low cost, high scalability, and automatic annotations, enabling the design of diverse scenarios, including rare or hazardous conditions. However, the simulation-to-reality gap limits their generalization capabilities. Real-world datasets, collected from actual environments, provide greater complexity and realism, reflecting sensor noise and dynamic interactions, making them essential for final model validation. Despite their reliability, they are costly and labor-intensive to acquire and annotate, often lacking scenario diversity. As a result, simulation datasets are preferred for algorithm development and pre-training, while real-world datasets remain critical for evaluation and deployment [36].
Table III: Comparison of Simulated and Real Datasets Characteristics
Characteristic Sim Real
Data Collection Cost $\checkmark$ $\times$×
Scene Diversity, Environmental Diversity $\checkmark$ $\times$×
Dynamic Interaction Complexity $\times$× $\checkmark$
Ability to Simulate New Scenes $\checkmark$ $\times$×
Coverage of Extreme/Dangerous Scenes $\checkmark$ $\times$×
Coverage of Long-Tail Scenarios $\checkmark$ $\times$×
Difficulty in Large-Scale Data Collection $\checkmark$ $\times$×
Processing Complexity $\times$× $\checkmark$
Data Labeling Accuracy $\checkmark$ $\times$×
Time Synchronization Precision $\checkmark$ $\times$×
Data Reality and Authenticity $\times$× $\checkmark$
Data Randomness and Noise $\times$× $\checkmark$
High Complexity and Diversity $\times$× $\checkmark$
Repeatability, Algorithm Iteration Speed, High Safety $\checkmark$ $\times$×
High Cost of Updates and Maintenance $\times$× $\checkmark$
Data Privacy Compliance $\checkmark$ $\times$×
Industry Certification and Recognition $\times$× $\checkmark$
III-A Simulated Datasets
III-A1 Simulation Platforms
Simulators are essential for CP research in AD, offering cost-effective, safe, and annotation-efficient alternatives to real-world data collection. They enable accurate generation of object attributes and automated annotations. Open-source platforms such as CARLA [99], SUMO [100], and AirSim [101] are widely used for data synthesis. CARLA supports diverse driving scenarios and multimodal sensors, including RGB, depth, radar, and LiDAR, but its LiDAR simulation has limitations in point cloud intensity and echo dropout modeling. SUMO provides vehicle and pedestrian trajectory simulation and can be co-simulated with CARLA to improve scene fidelity. The OpenCDA [102] framework integrates both platforms, offering standardized scenario management and benchmarks for perception and planning tasks.
III-A2 Simulation Datasets
Simulation datasets serve as vital testbeds for CP research, offering controllable environments, automated annotations, and support for diverse or rare scenarios. Representative datasets are built on CARLA, SUMO, or LiDARsim platforms. COMAP [79] and OPV2V [40] leverage CARLA-SUMO co-simulation for multi-agent perception across intersections and urban scenes, supporting detection, segmentation, and trajectory prediction. Extensions such as OPV2V+ [80] and OPV2V-H [83] improve agent diversity and simulate asynchronous sensing. DOLPHINS [88] and Adver-City [50] incorporate adverse weather to evaluate sensor robustness. V2X-Sim [41], Multi-V2X [89], and DeepAccident [51] expand support for high-resolution sensors, multi-agent scheduling, and accident prediction. These datasets provide critical foundations for benchmarking CP algorithms and exploring Sim2Real transfer, communication modeling, and generalization across scenarios.
III-B Real-World Datasets
Real-world CP datasets provide authentic sensor data and traffic interactions essential for AD research and model validation. DAIR-V2X [39] is the first large-scale multimodal real-world dataset focusing on V2I cooperation, while V2X-Seq [42] extends it with long-sequence data and traffic signal information. V2V4Real [53] is the first large-scale real-world multimodal V2V dataset, offering valuable support for Sim2Real studies despite limited CAV diversity. LUCOOP [81], collected with three CAVs using heterogeneous sensors, supports urban driving studies, while MARS [82] expands to four CAVs for improved stability and generalization. HoloVIC [85] and OTVIC [86] focus on V2I setups, with OTVIC simulating real-time vehicle-RSU communication to reflect dynamic transmission challenges in infrastructure-assisted perception.
IV Cooperative perception Scenario Datasets
CP datasets can be categorized according to different traffic environments to address the perception needs of AD systems in various scenarios. In this section, we classify them into three types: intersection, urban street, and highway scenarios. Fig. 5 shows the distribution of representative datasets across these categories. This chapter discusses the key datasets in each scenario and their characteristics, providing essential data support for CP research.
Figure 5: A Venn diagram showing the distribution of CP datasets across intersection, urban street, and highway scenarios. The overlapping areas indicate datasets applicable to multiple scenarios.
IV-A Intersection Datasets
Intersections represent one of the most complex traffic environments, where vehicles, pedestrians, and cyclists converge from multiple directions, leading to significant challenges such as occlusions and dynamic interactions. To enhance perception capabilities and environmental understanding, intersection CP datasets typically employ multi-sensor fusion and cross-view information integration techniques. The following datasets are among the most representative in this category.
Ko-PER [43] is one of the earliest datasets employing fixed LiDAR and cameras to monitor intersection dynamics. IPS300+ [74] improves spatial resolution with high-density 3D annotations, while LUMPI [76] introduces multi-view data under varying weather and lighting. TUMTraf-I [77] further targets annotation precision in adverse conditions, such as night and rain. HoloVIC [85] designs diverse holographic intersections to explore sensor placement strategies for CP. Several datasets incorporate intersection scenarios within broader V2X contexts. DAIR-V2X [39] supports V2I perception in urban crossings; DOLPHINS [88] includes intersection and T-junctions with weather variations; DeepAccident [51] focuses on accident prediction at intersections. V2XPnP [91] offers 163 intersection scenarios, supporting trajectory prediction and object tracking. These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion. However, challenges such as dynamic traffic variability, sensor blind zones remain. These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion.
IV-B Urban Street Datasets
Urban street CP datasets span V2V, V2I, and V2X paradigms, supporting tasks such as object detection, tracking, and trajectory prediction. CityFlow [44] offers multi-camera tracking data across arterials and residential areas. LUCOOP [81] collects time-synchronized multimodal data from three Connected Autonomous Vehicles (CAVs), including 3D point clouds, city models, and UWB ranging data. MARS [82] further extends to four CAVs under diverse weather, enhancing perception stability and agent collaboration. Several datasets strengthen roadside sensing capabilities. RCooper [60] provides the first large-scale real-world roadside dataset for multi-sensor infrastructure perception. RoScenes [78] delivers dense multi-view 3D data in complex urban traffic, supporting BEV perception. V2XPnP [91] captures diverse urban road scenes with multi-agent trajectories and detailed interactions. Together, these datasets address complex urban traffic environments, offering comprehensive data support for perception, fusion, and prediction in multi-agent V2X systems.
IV-C Highway Datasets
With the growing deployment of ICVs on highways, CP datasets for highway scenarios have become increasingly important for traffic management and AD safety. Highway environments demand long-range detection, lane-level perception, and robust environmental sensing, especially under high-speed conditions where precise synchronization is critical.
Although limited in number, highway CP datasets are gradually expanding. TUMTraf A9 [77] captures multi-weather data from Germany’s A9 highway for 3D object detection. OTVIC [86] introduces the first online-transmission, multi-view, multimodal dataset for V2I cooperative detection at 70–110 km/h. V2V4Real [53] includes 410 km of driving data, with 80% collected in highway environments. RoScenes [78] offers 14 ring-road scenarios with up to 800-meter perception range, enabling blind-spot-free coverage. H-V2X [52] combines radar and cameras for robust BEV perception in adverse weather, marking the first real-world, large-scale highway dataset of its kind. Other datasets such as DAIR-V2X [39], DOLPHINS [88], and CityFlow [44] also include partial highway scenarios, contributing to this domain.
V Datasets by Sensor Modality
Sensor modality diversity plays a pivotal role in enhancing the robustness of CP systems, particularly in complex and dynamic environments [103][104]. As illustrated in Fig. 6, typical sensors employed in CP include cameras, LiDAR, and Radar, each corresponding to distinct perception modalities such as RGB images, dense 3D point clouds, and sparse Radar point clouds. The respective strengths and weaknesses of these sensors are summarized in Table IV, providing a comprehensive overview of their capabilities under varying operational conditions. This section categorizes existing CP datasets based on their acquisition modalities, including image-only, LiDAR-only, and multimodal datasets.
Figure 6: Representative sensor types and corresponding perception modalities. (a) Camera and (b) RGB image; (c) LiDAR and (d) LiDAR point cloud; (e) 4D Radar and (f) sparse Radar point cloud. All data samples are derived from the V2X-Radar [57] dataset.
Image-only datasets, while cost-effective and semantically rich, are limited by poor depth estimation. LiDAR-only datasets offer precise 3D localization but are constrained by high sensor cost and limited semantic information. Multimodal datasets, especially those combining camera and LiDAR, dominate recent research due to their complementary strengths. Moreover, Radar-integrated datasets have gained attention for their resilience to adverse weather, marking a growing trend in sensor fusion for reliable perception [57].
Table IV: The Main Advantages and Disadvantages of Cameras, LiDAR, and Radar
Sensor Advantages Disadvantages
Camera • High resolution, capable of detecting color and texture information, good object classification ability.
• Low cost, mature technology.
• Recognizes traffic signs and lane markings.
• Relies on lighting conditions, reduced performance at night.
• Poor distance and speed measurement capability.
• Affected by weather conditions, performance degrades in adverse weather.
LiDAR • High-precision distance measurement, enables 3D environmental modeling.
• Not affected by lighting conditions, performs well at night.
• Wide field of view detection.
• Typically expensive.
• Performance degrades in rain, snow, and is affected by smoke, dust, and other weather conditions.
• Large data volume, complex processing.
Radar • Unaffected by weather and nighttime conditions, operates in all-weather environments.
• Accurate speed measurement, long detection range.
• Moderate cost, lower than LiDAR.
• Strong penetration (fog, rain, dust), strong anti-interference capability.
• Low resolution.
• Less sensitive to static objects, prone to false detections on metallic objects.
• Sparse detection points, weak reflection from pedestrians, making detection difficult.
• Cannot capture color or image information.
V-A Image-only Datasets
Image-only datasets offer high-resolution visual data, supporting tasks such as object detection and semantic segmentation. However, the absence of depth information limits their performance in occlusion handling, distance estimation, and robustness under varying lighting. CityFlow [44] spans large-scale urban scenarios, supporting cross-camera tracking and vehicle ReID, with extensive spatial and camera coverage. WIBAM [105], captured from infrastructure-mounted cameras, focuses on 2D/3D object detection. RoScenes [78], the largest multi-view roadside dataset to date, targets vision-based BEV perception, addressing challenges in multi-camera fusion and dense traffic scenes. Despite their effectiveness in visual tasks, image-only datasets have limited utility in high-precision localization and CP.
V-B LiDAR-only Datasets
LiDAR-only datasets provide accurate spatial geometry and are widely used for 3D detection, tracking, and map construction. While superior in depth perception, the absence of semantic cues limits their effectiveness in classification. V2V-Sim [47] and CODD [49] offer synthetic LiDAR data for perception and registration tasks. LUCOOP [81], a real-world dataset, supports multi-agent SLAM and urban navigation. Mixed Signals [92] enhances roadside cooperation through diverse LiDAR sensors. InScope [59] improves spatial awareness via distributed LiDAR deployments. Although point cloud datasets excel in 3D localization, their semantic limitations drive the shift toward multimodal fusion datasets in CP research.
V-C Multi-modal Datasets
Multi-modal datasets, particularly those fusing LiDAR and cameras, are central to CP research. By combining LiDAR’s precise 3D spatial sensing with camera-derived semantic richness, these datasets improve environmental understanding and robustness [106][107]. Recently, the integration of radar has further enhanced perception under adverse weather and for long-range detection. This section summarizes key LiDAR-Camera and LiDAR-Radar-Camera datasets.
LiDAR-Camera datasets include both infrastructure-based and cooperative datasets. Fixed-infrastructure datasets such as Ko-PER [43], IPS300+ [74], Rope3D [75], A9-Dataset [73], and LUMPI [76] use roadside LiDAR and cameras to produce accurate 3D annotations. H-V2X [52], RCooper [60], and TUMTraf-I [77] extend perception to challenging conditions such as rain and night. UAV and traffic camera datasets like INTERACTION [45] and BAAI-VANJEE [108] support interaction modeling and driving behavior analysis. CP datasets such as OPV2V [40], IRV2V [98], V2V4Real [53], LUCOOP [81], and MARS [82] collect multi-agent LiDAR-Camera data in both real and simulated settings. OPV2V-N [97] enhances camera robustness, while V2V-QA [84] introduces LLM-based collaborative tasks. CoPerception-UAV [35] supplements V2V data with aerial views. In V2X contexts, DAIR-V2X-C [39], V2X-Seq [42], HoloVIC [85], and OTVIC [86] support multi-modal detection, tracking, and planning. Simulated datasets such as V2X-Sim 2.0 [41], V2XSet [34], DOLPHINS [88], DeepAccident [51], V2X-Real [54], and Multi-V2X [89] further address diverse tasks from accident prediction to V2X data augmentation. LiDAR-Radar-Camera datasets such as T & J [46], OPV2V+ [80], OPV2V-H [83], and V2X-Radar [57] enhance CP under adverse conditions by leveraging radar’s robustness to weather and long-range sensing. In summary, LiDAR-Camera datasets remain the cornerstone of CP research, while LiDAR-Radar-Camera fusion introduces new capabilities for perception in challenging environments.
VI Perception Tasks and Metrics
With the development of CP, various datasets have been introduced to support key perception tasks [109][110] . This chapter categorizes them by task type, including object detection, semantic segmentation, object tracking, and trajectory prediction. Object detection focuses on 3D localization of targets, while semantic segmentation enables pixel-level scene understanding. Object tracking and trajectory prediction involve continuous monitoring and future motion inference of dynamic agents. For each task, we present representative datasets and their corresponding evaluation metrics to guide researchers in selecting appropriate datasets and assessment criteria.
VI-A Object Detection
Object detection is one of the most fundamental tasks in CP, encompassing both 2D and 3D object detection. In V2X scenarios, 3D object detection is particularly critical, often relying on LiDAR point clouds, camera images, or multi-modal data to achieve accurate detection of vehicles, pedestrians, and other objects in dynamic environments.
VI-A1 Datasets for Object Detection
Multiple datasets support V2X object detection across different collaboration modes and sensing modalities. DAIR-V2X-C [39] provides a real-world V2I benchmark for multi-modal detection. V2X-Sim 2.0 [41], offers 3D detection data for cooperative agents. HoloVIC [85] constructs holographic intersections for multi-view 3D sensing. V2X-Real [54] supports both V2I and V2V detection at scale, while V2V4Real [53] focuses on diverse road types for collaborative 3D detection. OPV2V [40] and OPV2V+ [80] simulate multi-agent cooperation, with the latter offering improved detection benchmarks. For roadside perception, RCooper [60] enhances infrastructure-based 3D detection under occlusions using multiple sensor layouts. Rope3D [75], while not a V2X dataset, provides monocular 3D detection benchmarks that benefit cooperative vision research. Overall, these datasets span real and simulated settings, support V2I and V2V collaboration, and offer diverse modalities, forming a solid foundation for V2X object detection studies.
VI-A2 Metrics for Object Detection
3D object detection in CP is primarily evaluated using Average Precision (AP)-based metrics and task-related performance indicators [111][112]. AP quantifies detection accuracy by computing precision at different Intersection-over-Union (IoU) thresholds [113]. Commonly used metrics include AP3D (based on 3D IoU) and APBEV (based on BEV IoU), as introduced by the KITTI benchmark [37].
Most CP datasets adopt AP or mAP as core metrics. RCooper [60] reports AP at multiple IoU thresholds. OPV2V [40] and OPV2V+ [80] evaluate vehicle detection at IoU@0.5 and IoU@0.7, with performance also measured within communication ranges. V2X-Real [54] employs both AP and mAP, using lower IoU thresholds (0.3/0.5) to account for varying object sizes. V2X-Sim 2.0 [41] focuses on BEV-based AP at multiple thresholds. Some datasets integrate communication cost into evaluation. V2V4Real [53] reports AP@0.5 and AP@0.7, and introduces Average MegaByte (AM) to quantify transmission overhead under synchronous and asynchronous conditions. DAIR-V2X [39] incorporates communication efficiency and limits evaluation to objects within the ego vehicle’s ROI. Rope3D [75] enhances AP-based evaluation with additional metrics such as Average Ground Center Similarity (ACS), Average Orientation Similarity (AOS), Average Area Similarity (AAS), and Average Four Ground Points Distance (AGD) to assess localization, orientation, and shape accuracy.
In addition to metric design, a range of benchmark methods have been evaluated across typical CP datasets. Table V summarizes the performance of representative algorithms on DAIR-V2X, V2V4Real, OPV2V, and V2XSet, offering insights into algorithm advancements and dataset challenges.
In summary, AP and IoU remain dominant evaluation standards in CP datasets, while extended metrics—such as BEV AP, communication cost, and asynchronous performance—offer deeper insights into real-world constraints. However, AP treats all missed detections equally, regardless of safety relevance. Future research should explore safety-aware evaluation frameworks tailored to CP tasks.
Table V: Performance Comparison of Representative Methods on Benchmark CP Datasets. All values are reported as AP@0.5 / AP@0.7. “-” indicates results not reported in the corresponding paper.
Method Year PUB Real-world Datasets Simulation Datasets
DAIR-V2X AP@0.5/0.7 Ref. V2V4Real AP@0.5/0.7 Ref. OPV2V AP@0.5/0.7 Ref. V2XSet AP@0.5/0.7 Ref.
F-Cooper [46] 2019 SEC 0.7370/0.5600 [96] 0.6930/0.4320 [96] 0.9070/0.8220 [97] 0.8750/0.7390 [97]
V2VNet [47] 2020 ECCV 0.7097/0.4763 [98] 0.6470/0.3360 [96] 0.8970/0.7400 [96] 0.8710/0.6460 [99]
V2VNetrobust [47] 2021 CoRL 0.6610/0.4860 [96] 0.5500/0.3090 [96] 0.9420/0.8540 [96] - -
V2X-Vit [34] 2022 ECCV 0.7398/0.6150 [102] 0.6590/0.4260 [104] 0.9460/0.8560 [103] 0.9100/0.8030 [99]
AttFuse [40] 2022 ICRA 0.7330/0.5530 [102] 0.7010/0.4540 [104] 0.9050/0.8160 [97] 0.8470/0.7400 [97]
Where2comm [35] 2022 NIPS 0.7520/0.5880 [96] 0.7040/0.4690 [96] 0.9440/0.8550 [96] 0.9130/0.8530 [100]
CoBEVT [114] 2022 CoRL 0.6390/0.5167 [102] 0.6480/0.4040 [104] 0.9330/0.8230 [104] 0.9033/0.8269 [102]
DiscoNet [115] 2021 NIPS 0.7370/0.5840 [96] 0.7360/0.4660 [96] 0.9160/0.7910 [103] 0.9078/0.8381 [102]
FPV-RCNN [116] 2022 RAL 0.6550/0.5050 [103] 0.7010/0.4790 [96] 0.8580/0.8400 [103] 0.8650/0.5630 [99]
CoAlign [117] 2023 ICRA 0.7460/0.6040 [103] 0.7090/0.4170 [96] 0.9660/0.9120 [103] 0.9190/0.8050 [96]
SCOPE [93] 2023 ICCV 0.6518/0.4989 [105] - - 0.8752/0.7505 [105] 0.8971/0.8062 [105]
V2VFormer [103] 2024 TIV - - 0.6540/0.3610 [107] 0.9170/0.8760 [107] - -
DI-V2X [118] 2024 AAAI 0.7882/0.6616 [98] - - - - 0.9270/0.8270 [98]
V2X-PC [119] 2024 arXiv 0.7689/0.6939 [101] - - - - 0.9283/0.8955 [101]
VI-B Semantic Segmentation
Semantic segmentation is a crucial task in CP, aiming to achieve pixel-level classification for enhanced environmental understanding. CP semantic segmentation datasets typically integrate LiDAR point clouds and camera images to enable finer scene interpretation, thereby improving AD systems’ perception in complex environments.
VI-B1 Datasets for Semantic Segmentation
Several datasets support semantic segmentation in CP tasks. V2X-Sim 2.0 [41] provides multi-agent LiDAR segmentation with dynamic obstacle, lane, and BEV semantic annotations. OPV2V [40] supports BEV semantic segmentation using LiDAR-camera fusion, while OPV2V+ [80] enhances cross-agent temporal alignment. OPV2V-N [97] focuses on improving image segmentation robustness. COMAP [79] offers BEV semantic segmentation in simulated urban scenarios. DeepAccident [51] provides accident scene segmentation with fine-grained semantic labels. DOLPHINS [88] supports V2X BEV segmentation under varied weather. Adver-City [50] collects segmentation data under extreme weather using dedicated semantic cameras.
Despite these advances, semantic segmentation in CP still faces several challenges, including inconsistent semantic representations across agents, and reduced robustness in adverse weather or sparse data conditions. Moreover, current datasets are limited in diversity and scale, especially for real-world multi-agent segmentation benchmarks. Future trends will emphasize the construction of large-scale real-world V2X segmentation datasets, cross-agent temporal consistency, and robustness under extreme conditions to support reliable and generalizable perception in complex collaborative scenarios.
VI-B2 Metrics for Semantic Segmentation
Semantic segmentation in CP tasks is commonly evaluated using Pixel Accuracy (PA), mean Pixel Accuracy (mPA), Intersection over Union (IoU), and mean IoU (mIoU) [120, 121]. PA measures the ratio of correctly classified pixels, mPA averages PA across classes to address imbalance, and mIoU reflects overall segmentation accuracy by averaging per-class IoU.
In CP scenarios, evaluation extends beyond single-agent performance to assess cross-agent consistency and modality alignment. V2X-Sim 2.0 [41] evaluates BEV semantic segmentation across categories such as lanes, vehicles, and buildings using mIoU. OPV2V-N [97] focuses on map-level segmentation (drivable area, lanes, vehicles), using IoU as the core metric. Effective CP segmentation evaluation requires traditional metrics and additional measures such as temporal stability, cross-modal IoU, and multi-agent coherence.
VI-C Tracking and Prediction
VI-C1 Datasets for Tracking and Prediction
Object tracking and trajectory prediction focus on modeling the temporal dynamics of moving agents and are essential for AD. CP improves tracking accuracy and trajectory prediction by enabling inter-agent information sharing, particularly under occlusion and complex traffic conditions.
Representative datasets include V2X-Seq [42], V2XPnP-Seq [91], and DAIR-V2X-Traj [56]. V2X-Seq offers long-sequence data for both V2V and V2I scenarios, while V2XPnP-Seq enhances temporal consistency and multi-view target association. DAIR-V2X-Traj focuses on V2I trajectory prediction with high-precision annotations. DeepAccident [51] addresses accident prediction by providing trajectory data in critical traffic scenarios.In urban-scale environments, CityFlow [44] supports cross-camera object tracking for traffic monitoring, and INTERACTION [45] provides detailed vehicle trajectories for interactive behavior modeling. RCooper [60] employs roadside sensors to support tracking in dense traffic. OPV2V [40] and V2V4Real [53] benchmark V2V tracking and trajectory prediction in simulated and real-world settings, respectively. V2X-Sim 2.0 [41] supports multi-agent tracking and prediction using BEV-based annotations.
Overall, CP datasets for tracking and prediction span V2V, V2I, and roadside sensing paradigms, providing diverse benchmarks for multi-agent temporal perception and forecasting in AD systems.
VI-C2 Metrics for Tracking and Prediction
Object tracking and trajectory prediction are central to CP, requiring precise evaluation of detection accuracy, tracking stability, and motion forecasting. Standard metrics have been developed to meet these requirements, with extensions introduced to address the complexities of multi-agent V2X scenarios.
For tracking tasks, the CLEAR MOT framework is widely adopted, incorporating Multiple Object Tracking Accuracy (MOTA), Precision (MOTP), and Identity F1 Score (IDF1). MOTA considers False Positives (FP), False Negatives (FN), and Identity Switches (IDSW), while MOTP measures average IoU between predicted and ground-truth trajectories. IDF1 evaluates consistency in identity association. Several datasets extend this framework. RCooper [60] and V2V4Real [53] incorporate Average MOTA (AMOTA), Average MOTP (AMOTP), and scaled AMOTA (sAMOTA) to assess tracking under varied object sizes and detection thresholds. DAIR-V2X-Traj [56] and V2X-Seq [42] also report MOTA, MOTP, and ID Switches for comprehensive analysis. To better reflect multi-agent collaboration, V2X-Sim 2.0 [41] introduces Higher Order Tracking Accuracy (HOTA), which jointly evaluates detection, association, and temporal stability. It also uses Byte Per Second (BPS) to quantify transmission cost, enabling analysis of the trade-off between bandwidth and accuracy.
Trajectory prediction tasks primarily use Average Displacement Error (ADE) and Final Displacement Error (FDE). ADE measures the mean offset over the entire predicted trajectory, while FDE focuses on the final point prediction accuracy. Datasets like INTERACTION [45] and V2X-Seq [42] adopt these metrics, with V2X-Seq further introducing Missing Rate (MR) to capture trajectory coverage completeness. In accident prediction, DeepAccident [51] proposes Volumetric Prediction Quality (VPQ) to jointly evaluate object detection and trajectory forecasting. It also introduces Accident Prediction Accuracy (APA), based on True/False Positives and Negatives, offering a safety-aware evaluation scheme for high-risk scenarios.
Overall, MOTA, MOTP, ADE, and FDE remain core metrics for tracking and prediction, while newer metrics like HOTA, BPS, VPQ, and APA enable more comprehensive evaluations tailored to V2X CP. These frameworks offer critical tools for benchmarking CP performance and guiding future system optimization.
VII Comprehensive Analysis of Collaborative Perception Datasets
To support robust and scalable development of CP systems, it is essential to comprehensively understand the characteristics, accessibility, and research impact of existing datasets [122]. This section systematically analyzes all reviewed CP datasets from three perspectives: their public availability and access mechanisms, the overall quality in terms of modality, annotation, and scene diversity, and the influence of representative datasets on the research community. These insights provide valuable references for dataset selection, algorithm benchmarking, and future dataset construction.
VII-A Availability and Access Methods of Datasets
Dataset accessibility plays a critical role in promoting research progress in CP [123] . Publicly available datasets reduce annotation costs and enable reproducibility, fostering innovation and collaboration. To assess the openness of current datasets, we categorize existing CP datasets based on their access methods.
First, a large portion of datasets can be directly downloaded from official websites, offering high accessibility without requiring registration or application. These include Ko-PER [43], CityFlow [44], LUMPI [76], TUMTraf-I [77], RoScenes [78], OPV2V [40], OPV2V+ [80], V2V4Real [53], LUCOOP [81], TUMTraf V2X [87], V2X-Sim 2.0 [41], DeepAccident [51], V2X-Real [54], Multi-V2X [89], Adver-City [50], MARS [82], Mixed Signals [92], and SCOPE [93]. In addition, some datasets are hosted on GitHub, often accompanied by source code, benchmark tools, or baseline implementations. Representative examples include V2V-Sim [47], OPV2V-H [83], CoPerception-UAV [35], CoopInf [48], V2XSet [34], and V2X-R [58], which facilitate rapid experimentation and method replication. Meanwhile, several datasets are distributed via cloud platforms such as Baidu Netdisk, Google Drive, or Alibaba Cloud. Access typically requires a request or registration process. These include H-V2X [52], DAIR-V2X-C [39], V2X-Seq [42], OTVIC [86], V2XReid [55], V2X-Radar [57], WHALES [90], and V2XPnP-Seq [91]. Furthermore, a number of datasets require formal application procedures, often involving submission of research proposals or institutional endorsements. Examples include BAAI-VANJEE [108], Rope3D [75], A9-Dataset [73], HoloVIC [85], V2V-QA [84], CODD [49], DAIR-V2X-Traj [56], INTERACTION [45], IPS300+ [74], DOLPHINS [88], InScope [59], and RCooper [60]. Finally, a few datasets do not publicly disclose their access procedures, possibly due to ongoing internal usage, limited availability, or institutional restrictions. These include WIBAM [105], T & J [46], IRV2V [98], OPV2V-N [97], CARTI [124], and V2X-DSI [125].
In summary, while many CP datasets are openly accessible via official websites, GitHub, or cloud storage, a significant portion remains gated behind application-based processes or lacks clear documentation. Improving data openness, unifying access procedures, and providing transparent distribution mechanisms will be essential to advancing CP research.
VII-B Comprehensive Quality Analysis of Datasets
Figure 7: Qualitative radar chart analysis of CP datasets. Each subplot represents a category of datasets and compares representative datasets across six key dimensions: Frame (number of frames captured in each dataset), Labels (annotation quality and category diversity), Agents (number and heterogeneity of collaborating entities), Scenarios (scene types and environmental coverage), Modalities (sensor types: camera, LiDAR, radar), and Tasks (supported perception tasks: object detection, tracking, segmentation, prediction). The analysis is qualitative and intended to provide an intuitive comparison of dataset characteristics, highlighting their strengths and focus areas in different CP contexts.
The quality of CP datasets directly impacts their applicability in real-world scenarios. As summarized in Table II and illustrated qualitatively in Fig. 7, we conduct a multi-dimensional analysis across six key aspects: data frame volume, annotation richness, agent diversity, scenario complexity, sensor modalities, and supported tasks. These dimensions reflect both dataset scalability and task coverage, providing a unified view for comparative assessment.
In terms of data volume, large-scale datasets such as RoScenes [78], INTERACTION [45], and DAIR-V2X-Traj [56] contain millions of frames or annotations, supporting long-sequence modeling and prediction tasks. For sensor diversity, OPV2V [40], V2X-R [58], and V2X-Radar [57] integrate camera, LiDAR, and radar modalities, enabling robust multimodal fusion. Regarding task support, datasets like V2X-Sim [41], DeepAccident [51], and SCOPE [93] offer multiple annotations for detection, tracking, segmentation, and prediction, enhancing their versatility for perception research. The radar charts in Fig. 7 further visualize dataset strengths and limitations across these dimensions within categories such as V2V, V2I, and roadside perception.
Overall, high-quality datasets exhibit complementary characteristics: real-world datasets like DAIR-V2X [39] and V2X-Real [54] emphasize fidelity and generalization, while simulation-based datasets like OPV2V+ [80] and Multi-V2X [89] excel in scale, flexibility, and controlled benchmarking. The increasing inclusion of diverse modalities, rich annotations, and comprehensive task coverage highlights a clear trend toward more holistic, multi-agent CP datasets. Future research would benefit from standardized benchmarking protocols across these high-quality datasets to promote reproducibility and cross-domain evaluation.
VII-C Summary of High-Influence Datasets
In the field of CP, several milestone datasets have gained prominence. This section summarizes key high-impact datasets across infrastructure-based, V2I, V2V, and V2X CP domains.
High-impact datasets are typically defined by high citation rates and widespread use in benchmark evaluations. For infrastructure-based perception, CityFlow [44], INTERACTION [45], and Rope3D [75] are widely used in multi-camera tracking and urban scene understanding. Representative methods such as BEVHeight [126], BEVHeight++ [127], and ViT-CoMer [128] have been developed using these datasets. In V2V perception, OPV2V [40] is one of the most cited simulation datasets, supporting algorithms like Where2comm [35], CoBEVT [114], and OPV2V-H [83]. In the V2I domain, DAIR-V2X [39] and its extensions (e.g., DAIR-V2X-Seq [42], DAIR-V2X-Traj [56]) provide large-scale real-world data with high-quality multimodal annotations, supporting tasks such as detection, tracking, and motion forecasting. These datasets underpin studies including CoCa3D [80], HM-ViT [129], BEV-V2X [130], and How2comm [131]. In the V2X category, V2X-Sim [41] supports BEV detection and semantic segmentation, and is frequently adopted by methods such as VoxFormer [132] and V2X-ViT [34].
In summary, datasets like CityFlow, INTERACTION, Rope3D, OPV2V, DAIR-V2X, and V2X-Sim have become cornerstones of CP research due to their broad adoption and benchmark availability. Newer datasets, including V2V4Real [53], Multi-V2X [89], TUMTraf V2X [87], and RCooper [60], are also gaining momentum and are expected to play increasingly important roles.
VIII Challenges and Frontiers
As research on CP continues to advance, datasets play a crucial role in facilitating multi-agent perception, V2X communication, and intelligent transportation systems. However, existing CP datasets still face numerous challenges, including insufficient data scale and diversity, limited availability, lack of standardization, domain adaptation issues, and concerns over privacy and security. Furthermore, with the rapid development of LLMs, their application in AD and CP datasets introduces new research trends. This chapter discusses these core challenges and explores potential future directions.
VIII-A Dataset Scale and Diversity
VIII-A1 Large-Scale Datasets
Large-scale datasets are essential for training deep learning-driven CP systems and improving their generalization ability. However, existing CP datasets are still significantly smaller than large-scale datasets in single-vehicle perception, such as the Waymo Open Dataset and nuScenes. While datasets like OPV2V, DAIR-V2X, and V2X-Sim provide relatively large sample sizes, they remain insufficient to comprehensively cover the complex and dynamic real-world traffic environment. Additionally, constructing large-scale datasets is costly and involves complex annotation processes, limiting their scalability. Future research may leverage automated annotation, synthetic data generation, and data augmentation techniques to reduce construction costs while enhancing dataset diversity and representativeness to better meet the needs of CP tasks.
VIII-A2 Datasets for Adverse Environment
Most current CP datasets primarily cover scenes under clear weather and good lighting conditions, with limited data available for adverse weather scenarios such as heavy rain, snow, fog, and extreme lighting. Adverse environments significantly impact sensor perception capabilities, such as LiDAR experiencing severe signal attenuation in rain and cameras suffering performance degradation under strong or low-light conditions. The lack of such data leads to poor generalization of existing perception models in extreme environments. Future research should focus on collecting data in adverse weather conditions and developing weather simulation techniques in virtual environments to supplement missing data. Additionally, leveraging multimodal sensor fusion and robustness-enhancing algorithms can improve model stability under challenging conditions.
VIII-A3 Datasets for Complex Scenarios
CP must handle complex scenarios such as high-density traffic, intersections, and lane changes. However, current datasets provide limited coverage of these scenarios. Object occlusion, dynamic interactions, and uncertainties in complex environments increase the difficulty of perception and prediction tasks. Future datasets should include diverse road types and driving behaviors to better simulate real-world driving conditions. Enhancing trajectory prediction and multi-agent interaction annotations will also provide a reliable benchmark for evaluating the performance of CP systems in complex scenarios.
VIII-A4 Datasets for Extreme Scenarios
Extreme scenarios are critical for evaluating the safety of AD systems. However, existing datasets contain few extreme scenario samples, limiting model generalization in these key situations. For example, the DeepAccident dataset attempts to provide traffic accident data, but its coverage of accident types remains incomplete. Future research should explore generating extreme scenario data through simulation, combined with real-world data augmentation, to enhance model responsiveness in hazardous situations. Moreover, anomaly detection and Out-Of-Distribution (OOD) analysis can help identify extreme scenarios, improving the safety and stability of CP systems.
VIII-A5 Geographic Diversity in Datasets
Traffic rules, driving behaviors, and road structures vary significantly across different countries and regions, posing challenges for deploying CP models in diverse geographical areas. Many publicly available datasets are region-specific; for instance, DAIR-V2X is primarily based on Chinese road data. The lack of geographic diversity may lead to performance degradation when models are deployed in different countries. Future datasets should incorporate traffic data from various regions to enhance geographic diversity and improve cross-region generalization capabilities. Additionally, transfer learning and domain adaptation techniques can help models better adapt to different traffic environments.
VIII-B Dataset Availability and Accessibility
Although multiple CP datasets are publicly available, some remain restricted, requiring special permissions or being limited to specific research collaborations. Furthermore, dataset storage and distribution methods are inconsistent, with some datasets hosted on unstable platforms, making access difficult for researchers. To promote research fairness and reproducibility, future efforts should focus on establishing more open data-sharing mechanisms, such as cloud-based platforms with unified data access interfaces. Additionally, standardizing dataset storage formats can improve portability and compatibility, allowing different research teams to efficiently utilize available data.
VIII-C Dataset Standardization and Benchmarking
Current datasets exhibit significant differences in annotation formats, category definitions, and evaluation metrics, making direct comparisons between studies challenging. For example, some datasets adopt the KITTI format for 3D object detection annotations, while others follow the nuScenes or Waymo formats. Additionally, inconsistencies in coordinate systems, units, and category definitions hinder data interoperability. Establishing a unified dataset standard that includes annotation specifications, data formats, and evaluation metrics will enhance compatibility and reproducibility, promoting standardized development in CP research.
VIII-D Domain Adaptation and Generalization
Domain adaptation remains a key challenge in improving dataset generalization. Due to differences in data collection environments, sensor configurations, and traffic scenarios, models trained on one dataset may not perform well on another. For instance, CP models trained in simulation environments often struggle when applied to real-world scenarios. Future research should explore unsupervised domain adaptation and self-supervised learning methods to enhance cross-domain generalization. Additionally, using Generative Adversarial Networks (GANs) to synthesize cross-domain data and leveraging large-scale pre-trained models for domain transfer learning could provide effective solutions.
VIII-E Privacy and Security
CP involves multi-agent data sharing, raising concerns about privacy protection and data security. Malicious attackers could intercept vehicle sensor data or manipulate transmitted data to disrupt collaborative decision-making. Furthermore, datasets may contain sensitive information, such as vehicle identification details and pedestrian facial features, increasing privacy risks. To address these issues, techniques such as differential privacy and federated learning can be employed to protect user privacy while maintaining data usability. Additionally, ensuring secure V2X communication mechanisms to prevent cyber attacks is an important direction for future research.
VIII-F LLMs and Datasets
Recently, LLMs have gained attention in the field of AD and intelligent transportation. LLMs can be used for automatic data annotation, synthetic data generation, and traffic scene analysis. However, their integration with V2X datasets is still in the early exploration stage. Key research directions include leveraging LLMs to enhance annotation efficiency and integrating LLMs for semantic understanding in multimodal perception tasks. Future trends may involve LLM-based data augmentation, AD knowledge graph construction, and the fusion of natural language with perception tasks to improve the intelligence level of V2X perception systems.
In summary, current CP datasets still face numerous challenges in terms of scale, availability, standardization, domain adaptation, and privacy security. Future directions include expanding dataset diversity and scale, establishing more comprehensive data-sharing mechanisms, promoting standardization efforts, and exploring the application of emerging technologies such as LLMs in dataset development. Advancements in these areas will further drive CP research and enhance the safety and reliability of AD systems.
IX Conclusion
This review presents a comprehensive review of CP datasets for AD, classifying them by collaboration paradigms, sensor modalities, data collection methods, environmental scenarios, and supported tasks. We analyze both real-world and simulation-based datasets across V2V, V2I, V2X, and I2I settings, evaluating their annotation quality, accessibility, scenario coverage, and benchmarking readiness. While existing datasets have significantly advanced CP research, limitations remain in data scale, coverage of rare and adverse conditions, agent diversity, and geographic generalization. To bridge these gaps, future datasets should incorporate long-tail scenarios such as traffic anomalies and extreme weather, include data from diverse regions, support heterogeneous agent types, and reflect realistic communication constraints. Establishing standardized annotation formats and unified evaluation protocols is essential for enabling fair comparisons and fostering cross-dataset generalization. Additionally, techniques such as automated labeling, high-fidelity simulation, and generative modeling—especially with the aid of large language models—offer promising directions for building scalable and semantically rich datasets. We hope this review provides valuable guidance for the design, construction, and benchmarking of future datasets, ultimately advancing CP systems in real-world AD applications.